please help me solve this equation i have no idea of what to do : 0.9(2x+8)=20-(x+5)
thanks !!

Found 2 solutions by stanbon, parthdesai27:
Answer by stanbon(75887)
Answer by parthdesai27(11)

18x+72 = 200-10x-50
=> 18x+72 = 200-10x-50
=> 28x = 150-72
=> x = 78/28 = 2.786































.